s Chapter 1 Determine the magnitude of the vector difflerence V v You will have to drow ลก oounteAgirefor thn graphical sohnun. V,:25 units Answers
Answer #1

(a) Magnitude of the difference of vector = 32.3 units

(b) Angle = 102.9 deg. from CC direction from + X-axis

For ve c h.. V.-4a-9 ニ옷 ;し , 116.6 す @5 Crs ts 엎 +25ト2s 3) uek = 그 (-3.52, +1? . o))-(23.01 +9.67) = +(-31.53 + 723) unh = 32-3 uuak